)I have a total of 112 yellow,brown and orange m&ms. if I have twice as many orange as yellow and 20 more yellow than brown how many of each color do I have?

Answer by checkley75(3666)

OR=2Y OR Y=OR/2
Y+20=B B=0R/2+20
Y+B+OR=112 THEN
OR/2+OR/2+20+OR=112
2OR=112-20
2OR=92
OR=92/2
OR=46 ORANGE M&MS
Y=46/2
Y=23 YELLOW M&MS
23+B+46=112
B=112-23-46
B=43 BROWN M&MS